# Approximate $2$-hop neighborhoods on incremental graphs: An efficient lazy approach

### v1

@@ 1. Introduction
In this paper, we consider the task of processing a possibly large, dynamic graph $G(V,E)$, incrementally provided as a stream of edge insertions, so that at any point of the stream it is possible to efficiently evaluate different queries that involve functions of the $h$-hop neighborhoods of its vertices. For a vertex $v\in V$, its $h$-hop neighborhood is simply the set of vertices that are within $h$ hops from $v$. In the remainder, $h$-hop neighborhoods are called $h$-balls for brevity. As concrete examples of query types we consider, one might want to estimate the size of the $2$-ball at a given vertex, or the Jaccard similarity between the $2$-balls centered at any given two vertices, or other indices of a similar flavor that depend on the intersection or union between $1$-balls and/or $2$-balls, just to mention a few.
Neighborhood-based indices are common in key mining tasks, such as link prediction in social (liben2003link, ) and biological networks (wang2023assessment, ) or to describe statistical properties of large social graphs (becchetti2008efficient, ). For example, $2$-hop neighborhoods are important in social network analysis and similarity-based link prediction (zhou2021experimental, ; zareie2020similarity, ; Sim-Nodes_Survey_2024, ), while accurate approximations of $h$-balls’ sizes are used to estimate key statistical properties of (very) large social networks (boldi2011hyperanf, ; backstrom2012four, ), or as link-based features in classifiers for Web spam detection (becchetti2008link, ).
@@
Nowadays, standard applications in social network analysis often entail dynamic scenarios in which input graphs evolve over time, under a sequence of edge insertions and possibly deletions (aggarwal2014evolutionary, ).
@@ 1.1. Our Contribution
In this paper, we propose an approach that trades some degree of accuracy for a substantial improvement in the average number of light updates. In a nutshell, upon an edge insertion, our algorithm performs the (two) corresponding heavy updates, but in general only a subset of the required light updates, according to a scheme that combines a threshold-based mechanism and a randomized, batch-update policy. Hence, for every vertex $u$, we only keep an approximation (a subset to be specific) of $u$’s $2$-ball. If $1$- and $2$-balls are represented with suitable data sketches, our approach affords constant average update cost per edge insertion. [^2] The particular sketch used depends on the neighborhood queries we want to be able to serve. When sketches are used, the cost of merging two neighborhoods corresponds to the cost of combining the corresponding sketches, which is typically a constant that depends on the desired approximation guarantees. For example, if we are interested in the Jaccard similarity between pairs of $1$- and/or $2$-balls, this cost will be proportional to the (constant) number of minhash values we use to represent each neighborhood. While the behavior and accuracy guarantees of most sketching techniques are well understood, the estimation error induced by lazy updates can be arbitrarily high in some cases. The main focus of this paper is on the latter aspect, which is absent in the static case but critical in the dynamic setting. Accordingly, we assume lossless representations of $1$-balls and approximate $2$-balls in our theoretical analyses in Sections 2 and 4, while we use standard sketching techniques to represent $1$- and $2$-balls in the actual implementations of the algorithms and baselines we consider in the experimental analysis discussed in Section 5.
@@
As positive as this result may sound, it begs the following questions from a careful reader: 1) Are the results above robust to adversarial sequences? 2) Is a performance analysis under random sequences representative of practical scenarios? More generally, does our lazy scheme offer significant practical advantages?
@@
As for the second question above, an analysis under random permutation sequences as the one in Section 3 is relatively common in the literature on dynamic edge streams and data streams (buriol2006counting, ; kapralov2014approximating, ; peng2018estimating, ; Hanauer22DynamicSurvey, ). Yet, one might rightly wonder about its practical significance for the task considered in this paper. We investigate this question in Section 5, where we conduct experiments on small, medium and large-sized, incremental graphs (whose main properties are summarized in Table 1). At least on the diverse sample of real networks we consider, experimental results on the estimation of key queries such as size and Jaccard similarity are consistent with the theoretical findings from Section 3. The main take-away is that, when using sketches to represent the $1$- and $2$-balls, the errors obtained with our lazy update policy are similar and fully comparable to those of the baseline, which performs all necessary light updates. At the same time, our algorithm proves to be significantly faster than the baseline, sometimes achieving a speedup of up to 90×90$\times$90 ×.
We finally remark that the datasets we consider are samples of real social networks. As such, they have relatively large local and global clustering coefficients [^4] At least the undirected ones. and thus low girth. Hence, our experimental analysis further supports the robustness of our theoretical findings: forcing our algorithm(s) into a worst-case behavior not only requires topologies characterized by a low girth, but also carefully crafted input sequences that are unlikely to occur in practice.
Remark.
@@
We finally stress that although we present them for the undirected case for ease of presentation and for the sake of brevity, our algorithms apply to directed graphs as well, [^5] Of course, in this case we have directed $h$-balls, i.e., sets of a vertices that can be reached in at most $h$ hops from a given vertex, or from which it is possible to reach the vertex under consideration in at most $h$ hops. while our analysis extends to the directed case with minor modifications.
@@ 1.2. Further related work
Efficient data structures for queries that involve $h$-balls of a dynamic graph turn out to be useful in different network applications. Besides those we mentioned earlier (becchetti2008link, ; zareie2020similarity, ; Sim-Nodes_Survey_2024, ), we cite here the work (cavallo20222, ), where the notion of 2-hop Neighbor Class Similarity (2NCS) is proposed: this is a new quantitative graph structural property that correlates with Graph Neural Networks (GNN) (scarselli2008graph, ; wu2022graph, ) performance more strongly and consistently than alternative metrics. 2NCS considers two-hop neighborhoods as a theoretically derived consequence of the two-step label propagation process governing GCN’s training-inference process.
@@ 2. Lazy-Update Algorithms
The dynamic (incremental) graph model we study can be defined as a sequence ${\mathcal{G}}=\{G^{(0)}(V,E^{(0)}),\ldots,$ $G^{(t)}(V,E^{(t)}),\ldots G^{(T)}(V,E^{(T)})\}$, where: (i) the set of vertices $V=\{1,\ldots,n\}$ is fixed, (ii) $T\leq\binom{n}{2}$ is the final graph, while (iii) $E^{(t)}$ is the subset of edges at time $t$. Note that this changes in every time step $t\geq 1$, as a new edge $e^{(t)}$ is inserted, so that $E^{(t+1)}=E^{(t)}\cup\{e^{(t)}\}$. We remark that our analysis and all our results can be easily adapted to a more general model that includes any combination of the following variants: (i) growing vertex sets, (ii) multiple insertions of the same edge, and (iii) directed edges (thus yielding directed graphs). However, the corresponding adaptations of our analysis would require significantly heavier notation and some technicalities that we decided to avoid for the sake of clarity and space.
@@
We next summarize notation that is extensively used in the remainder of the paper. For a vertex $v\in V$ of a graph $G(V,E)$, we define:
@@
The reader may have noticed that, in our notation above, the term $t$ does not appear: this is due to the fact that our analysis holds at any (arbitrarily-fixed) time step, which is always clear from context.
@@ 1 Function Insert($u,v$):
2       foreach $x\in\mathcal{N}(u)\cup\{u\}$ do
3             $B_{2}(v)\leftarrow B_{2}(v)\cup\{x\}$
4             $B_{2}(x)\leftarrow B_{2}(x)\cup\{v\}$
@@
7      foreach $x\in\mathcal{N}(v)\cup\{v\}$ do
8             $B_{2}(u)\leftarrow B_{2}(u)\cup\{x\}$
9             $B_{2}(x)\leftarrow B_{2}(x)\cup\{u\}$
@@
The magnitude of the changes (and the associated computational costs) induced by Insert$(u,v)$ vary. In particular, $B_{2}(u)$ can change significantly, as all vertices in $B_{1}(v)$ will be included in $B_{2}(u)$ (we refer to this as a heavy update). Instead, for any vertex $w\in\mathcal{N}(u)\setminus\{v\}$, $B_{2}(w)$ will grow by at most one element, namely $v$ (this is referred to as a light update). Symmetrically, the same holds for $v$ and for every $w\in\mathcal{N}(v)\setminus\{u\}$.
@@ 3. Random edge sequences
$\mathbb{P}\left(|\hat{B}_{2}(v)|<\frac{1-\alpha}{1+\varphi}|L_{2}(v)|\right)\leq e^{-\frac{2\alpha^{2}|L_{2}(v)|}{(1+\varphi)^{2}}}.$
@@ 4.2. Locally $\gamma$-sparse graphs
Let $\gamma\in\{0,1\ldots,n-1\}$. A graph $G(V,E)$ is said locally $\gamma$-sparse if for each vertex $u\in V$: (i) $\forall v\in L_{1}(u)$ the degree of $v$ in $G[L_{1}(u)]$ is at most $\gamma$, and (ii) $\forall w\in L_{2}(u)$ the degree of $w$ in $G[L_{1}(v)\cup\{w\}]$ is at most $\gamma+1$.
Observe that the class of locally $\gamma$-sparse graphs grows monotonically with $\gamma$, including all possible graphs for $\gamma=n-1$, while the most restricted class is obtained for $\gamma=0$. It is interesting to note that locally $\gamma$-sparse graphs are not necessarily sparse in absolute terms. For example, for $\gamma=0$, the class coincides with the well-known class of graphs with girth at least $5$: these graphs can have up to $\Theta(n^{\frac{3}{2}})$ edges assuming Erdös’ Girth Conjecture (erdos1965some, ) (the proof of such equivalence is given in Section A.2).
A first, preliminary analysis of our lazy-update approach considers the deterministic version of Algorithm 3, i.e., when $k=0$. It turns out that this version achieves an approximation ratio of $\frac{\gamma+1}{1-\varepsilon}$ and amortized cost $O(1/\varepsilon)$ (see Section A.2). So, the approximation accuracy decreases linearly in the parameter $\gamma$. Interestingly enough, we instead show that a suitable number of random light updates allows Algorithm 3 to perform much better than its deterministic version. This is the main result of this section and it is formalized in the next
@@
In this section, we use ^$\hat{\ell}_{v}$ and $\ell_{v}$ to denote the number of quasi-black and red edges, respectively, that connect $v$ to vertices in $L_{2}(u)$. Similarly, we use $b_{v}$ to represent the number of black edges of $v$ having the other endpoint in $L_{2}(u)$. Notice that ^$\hat{\ell}_{v}$ is a random variable that counts how many vertices out of $\ell_{v}$ are in $\hat{B}_{2}(u)$. We now proceed by first stating a property whose proof can be found in Appendix A.3.
@@ 5. Experimental analysis
The purpose of this experimental analysis, conducted on real datasets, is the validation of our lazy approach along three main axes: i) the qualitative consistency between the theoretical findings from Sections 2, 3 and 4 and the actual behavior of our lazy approach on real, incremental datasets; ii) the accuracy of our approach on two key neighborhood-based queries, namely, size and Jaccard similarity; iii) its computational savings with respect to non-lazy baselines on medium and large incremental graphs.
@@ 5.1. Experimental setup
Our experiments were performed on a machine with 2.3 GHz Intel Xeon Gold 5118 CPU with 24 cores, 192 GB of RAM, cache L1 32KB, shared L3 of 16MB and UMA architecture. The whole code is written in C++, compiled with GCC 10 and with the following compilation flags: -DARCH__\__X86__\__64 -Wall -Wextra -g -pg -O3 -lm.
@@
We considered real incremental graphs of different sizes, both directed and undirected, available from NetworkRepository (networkrepository, ). In our time analysis, we also extracted a large incremental dataset from a large static graph, namely soc-friendster (soc-friendster, ), available from SNAP (snapnets, ). Following previous work on dynamic graphs (chen2022dynamictree, ; Hanauer22DynamicSurvey, ), we generated an incremental graph by adding edges sequentially and in random order, starting from an empty graph. The main features of our datasets are summarized in Table 1.

### v2

@@ 1. Introduction
In this paper, we consider the task of processing a possibly large, dynamic graph $G(V,E)$, incrementally provided as a stream of edge insertions, so that at any point of the stream it is possible to efficiently evaluate different queries that involve functions of the $h$-hop neighborhoods of its vertices. For a vertex $v\in V$, its $h$-hop neighborhood is simply the set of vertices that are within $h$ hops from $v$. In the remainder, $h$-hop neighborhoods are called $h$-balls for brevity. As concrete examples of query types we consider, one might want to estimate the size of the $2$-ball at any given vertex, or the Jaccard similarity between the $2$-balls centered at any given two vertices, or other indices of a similar flavor that depend on the intersection or union between $1$-balls and/or $2$-balls, just to mention a few.
Neighborhood-based indices are common in key mining tasks, such as link prediction in social (liben2003link, ) and biological networks (wang2023assessment, ) or to describe statistical properties of large social graphs (becchetti2008efficient, ). For example, $2$-hop neighborhoods are important in social network analysis and similarity-based link prediction (zhou2021experimental, ; zareie2020similarity, ; Sim-Nodes_Survey_2024, ), while accurate approximations of $h$-balls’ sizes are used to estimate key statistical properties of (very) large social networks (boldi2011hyperanf, ; backstrom2012four, ), or as link-based features in classifiers for Web spam detection (becchetti2008link, ). Further tasks that may explicitly involve or benefit from $2$-hop neighborhood queries are accurate estimates of centrality measures that are widely adopted in social network analysis (freeman2002centrality, ; rochat2009closeness, ; bergamini2019computing, ) or mining of (usually large) bipartite networks that are frequent in certain applications (sariyuce2018peeling, ; pavlopoulos2018bipartite, ). We investigate the former example in detail in Section 6.
@@
Nowadays, standard applications in social network analysis often entail dynamic scenarios in which input graphs evolve over time, under a sequence of edge insertions and possibly deletions (aggarwal2014evolutionary, ). In many cases, such as co-authorship, citation or user-item association networks, the corresponding dynamic graphs are inherently or mostly incremental, i.e., either edges and/or vertices are added but not deleted, or insertions account for the vast majority of updates (leskovec2007graph, ; borner2004simultaneous, ). [^1] Of course, notable examples exist in which deletions are as important, such as some affiliation networks, or networks that describe on-line messaging behavior, where edges representing message exchanges become less and less important as they age.
@@ 1.1. Our Contribution
In this paper, we propose an approach that trades some degree of accuracy for a substantial improvement in the average number of light updates. In a nutshell, upon an edge insertion, our algorithm performs the (two) corresponding heavy updates, but in general only a subset of the required light updates, according to a scheme that combines a threshold-based mechanism and a randomized, batch-update policy. Hence, for every vertex $u$, we only keep an approximation (a subset to be specific) of $u$’s $2$-ball. If $1$- and $2$-balls are represented with suitable data sketches, our approach affords constant average update cost per edge insertion. [^3] The particular sketch used depends on the neighborhood queries we want to be able to serve. When sketches are used, the cost of merging two neighborhoods corresponds to the cost of combining the corresponding sketches, which is typically a constant that depends on the desired approximation guarantees. For example, if we are interested in the Jaccard similarity between pairs of $1$- and/or $2$-balls, this cost will be proportional to the (constant) number of minhash values we use to represent each neighborhood. While the behavior and accuracy guarantees of most sketching techniques are well understood, the estimation error induced by lazy updates can be arbitrarily high in some cases. The main focus of this paper is on the latter aspect, which is absent in the static case but critical in the dynamic setting. Accordingly, we assume lossless representations of $1$-balls and approximate $2$-balls in our theoretical analyses in Sections 2, 3 and 4, while we use standard sketching techniques to represent $1$- and $2$-balls in the actual implementations of the algorithms and baselines we consider in the experimental analysis discussed in Section 5.
@@
As positive as this result may sound, it begs the following questions from a careful reader: 1) Are the results above robust to adversarial sequences? 2) Is a performance analysis under random sequences representative of practical scenarios? 3) More generally, does our lazy scheme offer significant practical advantages?
@@
An analysis under random permutation sequences is relatively common in the literature on dynamic edge streams and data streams (buriol2006counting, ; kapralov2014approximating, ; peng2018estimating, ; Hanauer22DynamicSurvey, ). Yet, one might reasonably wonder about its practical significance for the task considered in this paper. We address these issues in Section 5, where we conduct experiments on small, medium and large-sized, incremental graphs (whose main properties are summarized in Table 1). At least on the diverse sample of real networks we consider, experimental results on the estimation of key queries such as size and Jaccard similarity are consistent with the theoretical findings from Section 3, while results on running times highlight the significance of our approach in practice: i) light updates crucially affect computational costs in practical scenarios; ii) when combined with sketches, our approach strikes an excellent balance between loss in accuracy and computational efficiency. In particular, its accuracy is comparable or only slightly worse than the baseline that performs all necessary light updates, while it achieves speed-ups that increase with network size and can be orders of magnitude with respect to the baseline on large or very large social networks (Tables 4 and 5). We finally note that our datasets are samples of real social networks. As such, they have relatively large local and global clustering coefficients [^5] At least the undirected ones. and thus low girth. Hence, our experimental analysis further supports the robustness of our theoretical findings: forcing our algorithm(s) into a worst-case behavior not only requires topologies characterized by a low girth, but also carefully crafted input sequences that are unlikely to occur in practice.
Use cases from real networks.
While neighborhood-based queries are extremely important in social network analysis, the main focus of this paper is on crucial, new challenges that arise in dynamic settings and on possible strategies to address them. For this reason, we investigate basic neighborhood queries without specific applications in mind. As simple as it may appear however, the incremental setting we consider provides a feasible framework to tackle important mining tasks over large, dynamic networks. As an example, we consider the problem of maintaining the harmonic centrality-based ranking of the vertices of an incremental graph (rochat2009closeness, ; boldi2014axioms, ), a key mining task in social network analysis (boldi2014axioms, ; freeman2002centrality, ). Exact solutions for this problem necessarily incur high computational costs in some cases, [^6] Even in the static setting (bergamini2019computing, ). while the study of dynamic strategies with update costs of practical interest is still an open area of research (HanauerHS22, ). In Section 6, we experimentally show that our approach is a very good fit for this task, affording continuous, accurate tracking of top-$k$ ranking vertices, thus highlighting its effectiveness for a key downstream task that would otherwise be hardly feasible on large, incremental graphs.
Many real networks are inherently bipartite, describing interactions between entities of two different types. Examples include many biological networks (pavlopoulos2018bipartite, ), social networks (LATAPY200831, ), user-URL or user-item interactions (leskovec2009meme, ) to name a few. Moreover, many of them are also inherently or mostly incremental in nature. [^7] Edge/vertex deletions are rare or absent. In such networks, the prediction of future interactions is a key mining task, for which (Jaccard) similarity between $2$-hop neighborhoods can prove a valuable index for prediction purposes. For example, in author-paper networks, the set $L_{2}(v)$ of vertices at distance exactly $2$ from an author $v$ is the set of $v$’s co-authors. Hence, for any two authors $u$ and $v$ with no currently co-authored paper, the Jaccard similarity between $L_{2}(u)$ and $L_{2}(v)$ can be used to estimate the probability of $u$ and $v$ co-authoring a paper in the next future. [^8] Our approach also allows estimation of the Jaccard similarity between $L_{2}(u)$ and $L_{2}(v)$.
Final remarks.
@@
We stress that although we present them for the undirected case for ease of presentation and for the sake of simplicity, our algorithms apply to directed graphs in general, [^9] Of course, in this case we have directed $h$-balls, i.e., sets of a vertices that can be reached in at most $h$ hops from a given vertex, or from which it is possible to reach the vertex under consideration in at most $h$ hops. while our analysis extends to the directed case with minor modifications. Moreover, our algorithms seamlessly address incremental settings in which new vertices may connect to the current graph, the latter being for example the typical scenario in co-authorship networks (borner2004simultaneous, ).
@@ 1.2. Further related work
Efficient data structures for queries that involve $h$-balls of a dynamic graph turn out to be useful in different network applications, as we mentioned earlier (becchetti2008link, ; zareie2020similarity, ; Sim-Nodes_Survey_2024, ).
@@ 2. Lazy-Update Algorithms
The dynamic (incremental) graph model we study can be defined as a sequence ${\mathcal{G}}=\{G^{(0)}(V,E^{(0)}),\ldots,$ $G^{(t)}(V,E^{(t)}),\ldots G^{(T)}(V,E^{(T)})\}$, where: (i) the set of vertices $V=\{1,\ldots,n\}$ is fixed, (ii) $T\leq\binom{n}{2}$ is the final graph, while (iii) $E^{(t)}$ is the subset of edges at time $t$. Note that this changes in every time step $t\geq 1$, as a new edge $e^{(t)}$ is inserted, so that $E^{(t+1)}=E^{(t)}\cup\{e^{(t)}\}$. We remark that our analysis and all our results can be easily adapted to a more general model that includes any combination of the following variants: (i) growing vertex sets, (ii) multiple insertions of the same edge, and (iii) directed edges (thus yielding directed graphs). However, the corresponding adaptations of our analysis would require significantly heavier notation and some technicalities that we decided to avoid for the sake of clarity.
@@
We next summarize notation that is used in the remainder of the paper. For a vertex $v\in V$ of a graph $G(V,E)$, we define:
@@
The reader may have noticed that, in our notation above, the term $t$ does not appear: this is due to the fact that our analysis holds at any (arbitrarily fixed) time step, which is always clear from context.
@@ 1 Function Insert($u,v$):
2       $B_{2}(u)\leftarrow B_{2}(u)\cup B_{1}(v)$
3       foreach $w\in\mathcal{N}(u)\setminus\{v\}$ do
4             $B_{2}(w)\leftarrow B_{2}(w)\cup\{v\}$
@@
7      $B_{2}(v)\leftarrow B_{2}(v)\cup B_{1}(u)$
8       foreach $w\in\mathcal{N}(v)\setminus\{u\}$ do
9             $B_{2}(w)\leftarrow B_{2}(w)\cup\{u\}$
@@
The magnitude of the changes (and the associated computational costs) induced by Insert$(u,v)$ vary. In particular, $B_{2}(u)$ can change significantly, as all vertices in $B_{1}(v)$ will be included in $B_{2}(u)$ (we refer to this as a heavy update, see Algorithm 1 at Algorithm 1). Instead, for any vertex $w\in\mathcal{N}(u)\setminus\{v\}$, $B_{2}(w)$ will grow by at most one element, namely $v$ (this is referred to as a light update, see Algorithm 1 at Algorithm 1). Symmetrically, the same holds for $v$ and for every $w\in\mathcal{N}(v)\setminus\{u\}$ (Algorithms 1 and 1).
@@ 3. Random edge sequences
(1) $\mathbb{P}\left(|\hat{B}_{2}(v)|<\frac{1-\alpha}{1+\varphi}|L_{2}(v)|\right)\leq e^{-\frac{2\alpha^{2}|L_{2}(v)|}{(1+\varphi)^{2}}}.$
@@ 4.2. Locally $\gamma$-sparse graphs
Let $\gamma\in\{0,1,\ldots,n-2\}$. A graph $G(V,E)$ is said locally $\gamma$-sparse if for each vertex $u\in V$: (i) $\forall v\in L_{1}(u)$ the degree of $v$ in $G[L_{1}(u)]$ is at most $\gamma$, and (ii) $\forall w\in L_{2}(u)$ the degree of $w$ in $G[L_{1}(v)\cup\{w\}]$ is at most $\gamma+1$.
Observe that the class of locally $\gamma$-sparse graphs grows monotonically with $\gamma$, including all possible graphs for $\gamma=n-2$, while the most restricted class is obtained for $\gamma=0$. It is interesting to note that locally $\gamma$-sparse graphs are not necessarily sparse in absolute terms. For example, for $\gamma=0$, the class coincides with the well-known class of graphs with girth at least $5$: these graphs can have up to $\Theta(n^{\frac{3}{2}})$ edges assuming Erdös’ Girth Conjecture (erdos1965some, ) (the proof of such equivalence is given in Section A.2).
A first, preliminary analysis of our lazy-update approach considers the deterministic version of Algorithm 3, i.e., when $k=0$. It turns out that this version achieves an approximation ratio of $\frac{\gamma+1}{1-\varepsilon}$ and amortized cost $O(1/\varepsilon)$ (see Section A.2). So, the approximation accuracy decreases linearly in the parameter $\gamma$. Interestingly enough, we instead show that a suitable number of random light updates allows Algorithm 3 to perform much better than its deterministic version. This is the main result of this section and it is formalized in the following theorem.
@@
In this section, we use ^$\hat{\ell}_{v}$ and $\ell_{v}$ to denote the number of quasi-black and red edges, respectively, that connect $v$ to vertices in $L_{2}(u)$. Similarly, we use $b_{v}$ to represent the number of black edges of $v$ having the other endpoint in $L_{2}(u)$. Notice that ^$\hat{\ell}_{v}$ is a random variable that counts how many vertices out of $\ell_{v}$ are in $\hat{B}_{2}(u)$. We now proceed by first stating a property, whose proof can be found in Section A.3.
@@ 5. Experimental analysis
The purpose of this experimental analysis, conducted on real datasets, is the validation of our lazy approach along three main axes: i) the qualitative consistency between the theoretical findings from Sections 2, 3 and 4 and the actual behavior of our lazy approach on real, incremental datasets; ii) the accuracy of our approach on two key neighborhood-based queries, namely, size and Jaccard similarity; iii) its computational savings with respect to non-lazy baselines on medium and large incremental graphs. An extensive and complete set of results is reported in Appendix B.
@@ 5.1. Experimental setup
Our experiments were performed on a machine with 2.3 GHz Intel Xeon Gold 5118 CPU with 24 cores, 192 GB of RAM, cache L1 32KB, shared L3 of 16MB and UMA architecture. The whole code is written in C++, compiled with GCC 10 and with the following compilation flags: -DARCH__\__X86__\__64 -Wall -Wextra -g -pg -O3 -lm. [^12] Code available at https://github.com/Gnumlab/graph_ball.
@@
We considered real incremental graphs of different sizes, both directed and undirected, available from NetworkRepository (networkrepository, ). In our time analysis, we also extracted a large incremental dataset from a large static graph, namely soc-friendster (soc-friendster, ), available from SNAP (snapnets, ). Following previous work on dynamic graphs (chen2022dynamictree, ; Hanauer22DynamicSurvey, ), we generated an incremental graph by adding edges sequentially and in random order, starting from an empty graph. The main features of these datasets are summarized in Table 1.